# Luminis Stellarum et Machina: Applications of Machine Learning in Light Curve Analysis

### v1

@@ 3.3 Semi-supervised Learning
Figure 8: Basic process of semi-supervised learning.
Figure 9: Overview of evaluation metrics. The figure is sourced from Rainio et al. (2024).
3.4 Evaluation Metrics
@@
Different evaluation metrics are used depending on the specific task, ensuring that models are assessed appropriately based on their objectives. Figure 9, taken from Rainio et al. (2024), provides a comprehensive overview of evaluation metrics for various ML scenarios, detailing the tasks they address, the values that must be computed for statistical testing, the potential questions these tests can answer, and the appropriate statistical tests for each case.
@@ 4.1 Transiting Exoplanet Detection
Table 2: Summary of machine learning methods applied to exoplanet detection datasets.
@@ 4.2 Variable Star Analysis
Beta type EB
W Ursae Majoris type EW
@@
W Virginis: period $<$8 d CWA
period $>$8 d CWB
@@
RRC type RRC
RRd type RRD
@@
Table 4: Summary of machine learning methods applied to variable star analysis.
@@
Kepler RF Long et al. (2012) introduced noisification to reduce survey-dependent feature mismatch.
@@ 4.3 Supernova classification
Early ML applications focused on engineered features derived from parametric LC fits or domain-specific metrics. The Supernova Photometric Classification Challenge (SNPCC) Kessler et al. (2010) was created as a standardized dataset to evaluate ML-based photometric classification methods for SN. Many subsequent studies tested their algorithms on the SNPCC dataset Newling et al. (2011); Richards et al. (2012); Karpenka et al. (2013); Gupta et al. (2016); Lochner et al. (2016); Charnock & Moss (2017); Ishida et al. (2019); Pasquet et al. (2019); Santos et al. (2020); de Oliveira et al. (2023).
DL revolutionized SN classification by allowing models to process raw flux measurements or minimally preprocessed LCs. Initial ML applications employed a variety of feature‐based techniques: Kernel Density Estimation (KDE) and Boosting methods Newling et al. (2011), Non‐linear Dimensionality Reduction with RFs Richards et al. (2012), ANNs Karpenka et al. (2013), Domain Adaptation with AL Gupta et al. (2016), and Naive Bayes, KNN, SVM, ANN, and BDTs Lochner et al. (2016).
@@
Recent classification pipelines increasingly incorporate generative models, Gaussian Process (GP) augmentation, and real-time alert systems. For example, the ParSNIP framework Boone (2021) utilizes Variational Autoencoders (VAEs) to perform classification. Similarly, Avocado Boone (2019) applies LightGBM combined with GP augmentation to classify transients photometrically. Real-time alert brokers like Fink Leoni et al. (2022) streamline the early identification of SNe from surveys like ZTF using AL strategies. CNN-based models such as SCONE Qu et al. (2021) apply 2D GP regression to multi-band LCs, while Photo-$z$SNthesis Qu & Sako (2023) uses CNNs to generate full redshift probability distributions. Temporal convolutional networks (TCNs) paired with LightGBM, as in the TLW model Li et al. (2024), further demonstrate the effectiveness of hybrid architectures for robust, survey-independent transient classification.
Table 6: Summary of machine learning methods applied to supernova classification.
@@
SNPCC Non-linear Dimension Reduction and RF Richards et al. (2012) proposed the non-linear dimension reduction technique to detect structure in a data base of SNe LCs.
@@
PLAsTiCC and Simulated SDSS-II SN data CNN Qu & Sako (2023) developed Photo-$z$SNthesis, a CNN-based method for predicting full redshift probability distributions from multi-band SNe LCs.
@@ 5 Challenges and Open Issues
The application of ML to LC analysis has transformed astronomical research, yet several significant challenges remain in building robust, interpretable, and survey-independent classification systems. These challenges are expected to intensify with the data influx from next-generation surveys such as the LSST, which is projected to detect tens of millions of transient events each night.
A central difficulty lies in the heterogeneity and sparsity of data. Many surveys, including ASAS-SN and Gaia, produce LCs with irregular sampling due to observational constraints such as weather or scheduling. This irregularity poses challenges for the direct application of DL models, which are typically designed for regularly sampled data. Furthermore, models trained on data from one survey, such as Kepler, often fail to generalize to others like TESS, due to differences in cadence, noise properties, and photometric filters. The problem is compounded by the scarcity of high-quality labels, as spectroscopic confirmations are limited, especially for rare transient classes such as SLSNe.
Another major challenge involves the interpretability of ML models and their consistency with physical principles. ML models often function as black boxes, providing predictions without clear explanations. This lack of transparency limits their utility in cosmological studies that require rigorous uncertainty quantification and interpretability. There is a growing interest in hybrid approaches that combine data-driven learning with physics-based priors.
Real-time processing and early classification are also critical challenges. LSST’s alert stream will require sub-minute response times, which strain the capabilities of even highly optimized neural networks. Moreover, many existing classification techniques rely on full-phase LCs for accurate classification. In contrast, real-time systems must operate on partial data, often limited to early stages such as the rising or plateau phases of the LC. Developing models that can provide reliable early-time classifications is a key area of ongoing research.
Lastly, class imbalance and anomaly detection present persistent obstacles. Rare events such as kilonovae or luminious red novae are difficult to detect using standard classification methods. Traditional oversampling techniques are often insufficient for handling such imbalance, and alternative generative or data augmentation strategies are needed. In addition, most ML classifiers operate under a closed-set assumption, recognizing only a fixed set of known classes. However, LSST and similar surveys are expected to discover entirely new types of transients.

### v2

@@ 4.3 Semi-supervised Learning
[…]
Figure 9: Basic process of SSL. Cylinders represent databases, rectangles indicate data processing steps, and the rounded rectangle shows the final output.
4.4 Uncertainty and Bias
In ML models, uncertainty arises from two primary sources. Epistemic uncertainty (model uncertainty) reflects the lack of knowledge due to limited or unrepresentative training data and can be reduced by collecting additional observations or using ensemble or Bayesian methods. For example, a CNN trained on a limited set of supernova LCs may be uncertain when classifying rare or novel transients. Aleatoric uncertainty (data uncertainty) originates from inherent noise in measurements, such as photometric errors or instrumental noise, and cannot be reduced by more data. Bias can also occur systematically due to model assumptions or dataset composition; for instance, overrepresentation of bright galaxies may lead to misclassification of faint objects. Characterizing both uncertainty and bias is crucial in astronomy, as it informs the reliability of classifications, the confidence in parameter estimates, and the prioritization of follow-up observations. Techniques such as Monte Carlo dropout, Bayesian neural networks, and careful data augmentation are employed to quantify and mitigate these effects across the models.
4.5 Evaluation Metrics and Validation
4.5.1 Evaluation Metrics
@@
In regression problems, where the target variable is continuous, model performance is commonly evaluated using metrics such as the Mean Absolute Error (MAE) and the Mean Squared Error (MSE). The MAE is defined as:
$\text{MAE}=\frac{1}{n}\sum_{i=1}^{n}|y_{i}-\hat{y}_{i}|,$
where $n$ is the number of instances in the validation or test set, $y_{i}$ is the true value, and $\hat{y}_{i}$ is the model’s prediction. The MSE is given by:
$\text{MSE}=\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}.$
Moreover, additional evaluation metrics are also found in the literature, such as the Normalized Median Absolute Deviation, the Continuous Ranked Probability Score, and the Probability Integral Transform (see e.g., D’Isanto & Polsterer (2018); D’Isanto et al. (2018)).
In classification tasks, where the target variable is discrete, commonly used evaluation metrics include Classification Accuracy, the Confusion Matrix, and the Area Under the ROC Curve (AUC).
Classification accuracy is defined as the ratio of correctly predicted labels to the total number of predictions:
$\text{Accuracy}=\frac{\text{Number of correct predictions}}{\text{Total predictions}},$
with values ranging between 0 and 1. It is most appropriate when classes are balanced and misclassifications are equally important.
The confusion matrix provides a more detailed view of performance, especially in multi-class settings. Figure 10 shows an example of a confusion matrix, taken from Akhmetali et al. (2024), who trained a CNN model to distinguish between 5 classes of variable stars. Each row corresponds to the true class, and each column to the predicted class. Ideally, the matrix is diagonal, indicating perfect classification. Normalizing the matrix removes dependence on class size, allowing for easier comparison across categories.
Figure 10: Example of a confusion matrix taken from Akhmetali et al. (2024), who trained a CNN model to distinguish between 5 classes of variable stars. The confusion matrix displays the percentage distribution of true class labels against the predicted class labels. In the best-case scenario, the confusion matrix is purely diagonal, with all off-diagonal elements equal to zero, indicating that all instances were correctly classified.
The Receiver Operating Characteristic (ROC) curve is a standard diagnostic tool for binary classification. It plots the True Positive Rate (TPR) against the False Positive Rate (FPR):
$\text{TPR}=\frac{\text{True Positives}}{\text{Total Actual Positives}},$
$\text{FPR}=\frac{\text{False Positives}}{\text{Total Actual Negatives}}.$
By varying the classification threshold, the ROC curve visualizes model performance across different operating points. The Area Under the Curve (AUC) provides a summary measure of overall performance, with higher AUC values indicating better discrimination. ROC curves are valuable for comparing models and for guiding the selection of classification thresholds that balance sensitivity and specificity, as illustrated in Figure 11.
In multiclass settings, ROC analysis requires adaptation because TPR and FPR are inherently defined for binary outcomes. One approach is the one-vs-rest (OvR) strategy, in which each class is treated as the positive class while all other classes are treated as negative, producing a separate ROC curve for each class. Alternatively, macro-averaging computes the AUC for each class and averages them equally, giving each class the same weight, while micro-averaging aggregates true positives, false positives, and false negatives across all classes before calculating TPR and FPR, effectively weighting classes by prevalence. Interpreting ROC results in multiclass problems can be challenging, particularly when dealing with imbalanced datasets or more than three classes, and visualizing multiple ROC curves simultaneously may be difficult.
As an alternative to ROC, Precision-Recall (PR) curves are often more informative for imbalanced datasets because they emphasize performance on the positive class. PR curves can also guide threshold selection by highlighting points that optimize the trade-off between precision and recall.
By analyzing ROC or PR curves, it is possible to choose classification thresholds that optimize specific metrics such as the F1 score, balanced accuracy, or sensitivity, rather than relying on a default threshold of 0.5. This is especially important in astronomy, where rare events, such as supernovae or transient detections, require prioritizing sensitivity to avoid missing important signals.
Figure 11: An illustration of a ROC curve taken from Baron (2019), which plots the TPR against the FPR. Ideally, a model should achieve a TPR of 1 with a FPR of 0. The black dashed line represents the performance of a random classifier. By adjusting the model’s hyperparameters, different points on the ROC curve are generated using validation set results. The pink curve corresponds to model 1, with points A, B, and C indicating specific hyperparameter settings. The purple curve shows the performance of model 2. The area under the ROC curve (AUC) provides a measure for model comparison - model 1 performs better in this case.
4.5.2 Validation Methods
To reliably assess the generalization ability of ML models, it is crucial to evaluate them on data that was not used during training. This is achieved through various validation methods, which help prevent overfitting and provide unbiased estimates of model performance.
Holdout Validation (HV) Kim (2009) is the most common approach for evaluating ML models. In this approach, the dataset is split into training, validation, and test sets. A typical split uses 70% for training, 15% for validation, and 15% for testing, though these ratios may vary depending on dataset size and variability. For small datasets, a larger training proportion is preferred to ensure sufficient learning. As datasets grow, smaller validation sets can still capture data diversity effectively. The training set is used to fit model parameters, while the validation set is used for hyperparameter tuning—a process that, although computationally intensive, can be automated and parallelized. The final model is evaluated on the unseen test set. HV is used in various astronomical tasks, such as classification of spectral data Solorio-Ram´ırez et al. (2023), estimation of galaxy redshifts Ball et al. (2008), classification of sterllar variability Audenaert et al. (2021), and predicting stellar rotations Gomes et al. (2024).
Cross-validation (CV) Stone (1974) is a resampling technique used to obtain an unbiased estimate of a model’s performance. Compared to simple holdout validation, it generally offers a more reliable assessment of a model’s generalization error, particularly when working with small datasets. Several cross-validation strategies commonly used for model evaluation are outlined below. CV is used in astronomy for estimation of photometric redshifts of quasars Ball et al. (2007), Star–Galaxy–QSO classification Bai et al. (2018), calibration of ages for the APOGEE DR17 catalogue Boulet (2024), and classification of variabile sources Richards et al. (2012b).
K-Fold Cross-Validation (KFCV) Bengio & Grandvalet (2004) partitions the dataset randomly into $k$ disjoint subsets, or folds. In each iteration, one fold is held out as the validation set, while the remaining k1k-1 folds are used for training. This process repeats $k$ times, ensuring each fold serves exactly once as the validation set. The average of the performance metrics across all iterations provides a robust estimate of the model’s validation error. KFCV is used in various works, such as classification of variable stars Akhmetali et al. (2024), star–galaxy separation Machado et al. (2016b), protostellar classification Miettinen (2018), and identification and fitting eclipse maps of exoplanets Hammond et al. (2024).
Stratified K-Fold Cross-Validation (SKFCV) Prusty et al. (2022) is particularly useful in classification tasks with imbalanced class distributions. Unlike standard KFCV, stratified KFCV preserves the original class proportions within each fold. This ensures that minority classes are adequately represented, yielding more stable and representative performance estimates. SKFCV is used for prediction of habitability of exoplanets Raminaidu et al. (2023), classification of massive stars in nearby galaxies Maravelias et al. (2022), spectral classification of OB stars Kyritsis et al. (2022), and detection of quasi-periodic oscillations in X-ray binaries Kiker et al. (2023).
Leave-One-Out Cross-Validation (LOOCV) Wong (2015) is a special case of KFCV where $k=n$, the total number of data points. Each iteration uses a single sample for validation and the remaining n1n-1 samples for training. LOOCV provides nearly unbiased performance estimates, making it well-suited for very small or imbalanced datasets. However, it is computationally intensive and often impractical for large datasets or models with long training times. SKFCV is for exoplanet atmospheric analysis Welbanks et al. (2023), selection of optically variable active galactic nuclei De Cicco et al. (2025), analysis of stellar LCs Pan et al. (2024), and exoplanet detection Pimentel et al. (2024).
Nested Cross-Validation (NCV) Varma & Simon (2006) addresses the challenge of reliable hyperparameter tuning and unbiased performance estimation. The data is split into outer and inner loops of cross-validation. The inner loop is used for model selection (e.g., tuning hyperparameters), while the outer loop assesses the generalization error. This method helps avoid optimistic bias that can occur when hyperparameters are tuned and evaluated on the same data. NCV is used for star/galaxy classification Chao et al. (2020), cosmological inference from non-linear scales Yuan et al. (2024), classification of variable stars Long et al. (2012a), and prediction of galaxy metallicity from three-colour images Wu & Boada (2019).
Validation methods not only influence model selection but also help assess uncertainty, bias, and variance. For instance, large variations in performance across folds may indicate that the model is sensitive to the training data and prone to overfitting. Similarly, if the model consistently underperforms across validation sets, it may be underfitting or affected by biased assumptions.
Despite the common use of performance metrics and validation schemes, explicit statistical comparisons between models are still uncommon in astronomy. This is problematic because reported differences in metrics (e.g., accuracy or AUC) may not be statistically significant, especially with noisy or imbalanced data. Statistical significance tests help assess whether performance improvements are robust rather than due to random fluctuations Demšar (2006). For instance, McNemar’s test McNemar (1947) evaluates paired classification outcomes, while paired $t$-tests or Wilcoxon signed-rank tests Higgins (2004) can compare cross-validation scores. Bootstrap and permutation methods can also provide confidence intervals for performance differences Good (2005). Incorporating such tests offers a more rigorous foundation for model selection and helps prevent overstatement of results.
@@ 5.1 Transiting Exoplanet Detection
Table 2: Summary of ML methods applied to exoplanet detection datasets.
@@ 5.2 Variable Star Analysis
Eclipsing binary: Beta type EB
Eclipsing binary: W Ursae Majoris type EW
@@
W Virginis: period <<8 d CWA
W Virginis: period >>8 d CWB
@@
RR Lyrae: RRC type RRC
RR Lyrae: RRd type RRD
@@
Table 4: Summary of ML methods applied to variable star analysis.
@@
Kepler RF Long et al. (2012b) introduced noisification to reduce survey-dependent feature mismatch.
@@ 5.3 Supernova classification
Early ML applications focused on engineered features derived from parametric LC fits or domain-specific metrics. The Supernova Photometric Classification Challenge (SNPCC) Kessler et al. (2010) was created as a standardized dataset to evaluate ML-based photometric classification methods for SN. Many subsequent studies tested their algorithms on the SNPCC dataset Newling et al. (2011); Richards et al. (2012a); Karpenka et al. (2013); Gupta et al. (2016); Lochner et al. (2016); Charnock & Moss (2017); Ishida et al. (2019); Pasquet et al. (2019); Santos et al. (2020); de Oliveira et al. (2023).
DL revolutionized SN classification by allowing models to process raw flux measurements or minimally preprocessed LCs. Initial ML applications employed a variety of feature‐based techniques: Kernel Density Estimation (KDE) and Boosting methods Newling et al. (2011), Non‐linear Dimensionality Reduction with RFs Richards et al. (2012a), ANNs Karpenka et al. (2013), Domain Adaptation with AL Gupta et al. (2016), and Naive Bayes, KNN, SVM, ANN, and BDTs Lochner et al. (2016).
@@
Recent classification pipelines increasingly incorporate generative models, Gaussian Process (GP) augmentation, and real-time alert systems Martínez-Palomera et al. (2022); Boone (2021); Smith & Geach (2023). For example, the ParSNIP framework Boone (2021) utilizes Variational Autoencoders (VAEs) to perform classification. Similarly, Avocado Boone (2019) applies LightGBM combined with GP augmentation to classify transients photometrically. Real-time alert brokers like Fink Leoni et al. (2022) streamline the early identification of SNe from surveys like ZTF using AL strategies. CNN-based models such as SCONE Qu et al. (2021) apply 2D GP regression to multi-band LCs, while Photo-zzSNthesis Qu & Sako (2023) uses CNNs to generate full redshift probability distributions. Temporal convolutional networks (TCNs) paired with LightGBM, as in the TLW model Li et al. (2024), further demonstrate the effectiveness of hybrid architectures for robust, survey-independent transient classification.
Table 6: Summary of ML methods applied to supernova classification.
@@
SNPCC Non-linear Dimension Reduction and RF Richards et al. (2012a) proposed the non-linear dimension reduction technique to detect structure in a data base of SNe LCs.
@@
PLAsTiCC and Simulated SDSS-II SN data CNN Qu & Sako (2023) developed Photo-zzSNthesis, a CNN-based method for predicting full redshift probability distributions from multi-band SNe LCs.
@@ 6 Challenges and Open Issues
The application of ML to LC analysis has revolutionized astronomical research, yet significant challenges remain in developing robust, interpretable, and survey-independent classification systems. These challenges are expected to intensify with upcoming data-intensive surveys such as the LSST, which will detect tens of millions of transient events nightly.
A key difficulty lies in the heterogeneity and sparsity of data. Surveys like ASAS-SN and Gaia produce irregularly sampled LCs due to observational constraints such as weather or cadence gaps. This irregularity complicates the use of DL models, which typically assume regularly sampled inputs. Moreover, models trained on one survey (e.g., Kepler) often perform poorly when applied to others (e.g., TESS), due to differences in cadence, noise levels, and photometric filters Valizadegan et al. (2022). These issues are further exacerbated by limited spectroscopic follow-up, especially for rare classes such as SLSNe.
Model interpretability and physical consistency are additional concerns. Many ML models act as black boxes, offering little insight into their predictions. This lack of transparency hinders their adoption in cosmological studies that require rigorous uncertainty quantification and interpretability. Hybrid approaches that incorporate physical priors into data-driven models are a promising direction.
Real-time processing and early classification present another critical challenge. LSST will require sub-minute response times for alert streams, straining even highly optimized models. Furthermore, most classifiers rely on full-phase LCs, whereas early-time classification should operate with incomplete data—often limited to the rising or plateau phases. Building models that can deliver accurate early-stage predictions is an active area of research.
Class imbalance and anomaly detection remain persistent issues. Rare transients such as kilonovae and luminous red novae are difficult to detect using conventional techniques. Oversampling strategies often fall short, prompting interest in generative and data augmentation methods. Moreover, most classifiers assume a closed-set of known classes, yet upcoming surveys will likely discover entirely new types of transients.
Generative models—such as GANs, VAEs, and diffusion models—are emerging as powerful tools for addressing these challenges. They enable synthetic augmentation of rare classes, imputation of missing or irregular data, and cross-survey domain adaptation. While still in early development, generative approaches show strong potential to improve the robustness and generalizability of ML-based LC pipelines.